OpenStudy (anonymous):

Solve the quadratic k^2-4k+13=0

OpenStudy (jadeishere):

Okay, you could use the quadratic formula! a = 1 (k^2) b = -4 (-4k) c = 13 So \[x = \frac{-b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

use this formula

OpenStudy (anonymous):

and put value and get your ans

OpenStudy (anonymous):

i got  which was wrong

OpenStudy (jadeishere):

plug in the values for the formula -b = 4 since -(-4) = 4 b^2 = 16 4(a)(c) = 4 * 1 * 13 = 54 2a = 2 so with that \[x = \frac{ 4 \pm \sqrt{16 - 54} }{ }\]

OpenStudy (anonymous):

5 and -1

OpenStudy (jadeishere):

over 2

mathslover (mathslover):

4*1*13 = 52 (correction)

OpenStudy (jadeishere):

Nono, help her get the answer don't just say it

OpenStudy (jadeishere):

Oh, oops /)_- thanks mathslover

mathslover (mathslover):

You're welcome.

OpenStudy (anonymous):

k got it so x=-14,22

OpenStudy (jadeishere):

let me work it out completely

OpenStudy (jadeishere):

Will you tell me how you got -14 and 22? I got different answers

OpenStudy (anonymous):



OpenStudy (jadeishere):

Okay, so your mistake was at x = -36 / 2 the sqrt sign is looking for a value that multiplied by itself = 36. so sqrt(36) = 6

OpenStudy (twopointinfinity):

... congrats you get \[ \begin{array}l\color{red}{\text{F}}\color{orange}{\text{r}}\color{grey}{\text{e}}\color{green}{\text{e}}\color{blue}{\text{-}}\color{purple}{\text{C}}\color{purple}{\text{u}}\color{red}{\text{p}}\color{orange}{\text{c}}\color{grey}{\text{a}}\color{green}{\text{k}}\color{blue}{\text{e}}\color{purple}{\text{s}}\color{purple}{\text{}}\end{array} \]

OpenStudy (anonymous):

oops i forgot it was still square root! thanks